# Decoupling Contrastive Decoding: Robust Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ Abstract
Although multimodal large language models (MLLMs) exhibit remarkable reasoning capabilities on complex multimodal understanding tasks, they still suffer from the notorious “hallucination” issue: generating outputs misaligned with obvious visual or factual evidence. Currently, training-based solutions, like direct preference optimization (DPO), leverage paired preference data to suppress hallucinations. However, they risk sacrificing general reasoning capabilities due to the likelihood displacement. Meanwhile, training-free solutions, like contrastive decoding, achieve this goal by subtracting the estimated hallucination pattern from a distorted input. Yet, these handcrafted perturbations (e.g., add noise to images) may poorly capture authentic hallucination patterns. To avoid these weaknesses of existing methods, and realize “robust” hallucination mitigation (i.e., maintaining general reasoning performance), we propose a novel framework: Decoupling Contrastive Decoding (DCD). Specifically, DCD decouples the learning of positive and negative samples in preference datasets, and trains separate positive and negative image projections within the MLLM. The negative projection implicitly models real hallucination patterns, which enables vision-aware negative images in the contrastive decoding inference stage. Our DCD alleviates likelihood displacement by avoiding pairwise optimization and generalizes robustly without handcrafted degradation. Extensive ablations across hallucination benchmarks and general reasoning tasks demonstrate the effectiveness of DCD, i.e., it matches DPO’s hallucination suppression while preserving general capabilities and outperforms the handcrafted contrastive decoding methods. Code will be released.
@@ 1 Introduction
Today’s multimodal large language models (MLLMs) liu2024improved ; chen2023minigpt ; bai2023qwen ; dubey2024llama ; chen2024expanding have demonstrated remarkable general reasoning capabilities by integrating visual and textual understanding, facilitating applications such as medical image analysis li2024llava ; liu2023medical and multimodal search engines caffagni2024wiki . Despite their versatility, a critical limitation persists: MLLMs may generate outputs that contradict obvious factual evidence or misrepresent visual inputs, known as the hallucination problem bai2024hallucination ; liu2024survey . For instance, models may describe objects absent from an image (e.g., claiming a “dog” in a cat-only scene) or fabricate implausible relationships (e.g., asserting “a person riding a bicycle” when only a bicycle is present). Such hallucinations erode users trust and hinder deployment in high-stakes domains like healthcare li2024llava or autonomous driving cui2024survey .
Figure 1: Comparison between hallucination mitigation methods and our proposed DCD. (a) Training-based method: DPO rafailov2024direct . $v$, $x$, $y^{+}$, and $y^{-}$ stand for images, questions, position responses, and negative responses in preference datasets, respectively. $\theta$ denotes the parameter of the model. $\alpha$ is the coefficient in contrastive decoding. (b) Training-free method: VCD leng2024mitigating . $v^{+}$ and $v^{-}$ are positive and negative visual inputs for MLLM in the inference stage.
@@
    We first propose a way to decouple the positive and negative sample learning in preference datasets, which tries to achieve robust hallucination mitigation in MLLMs.
@@
    Our method even can learn from negative samples solely. We reveal that negative samples are more important than positive samples in the RLHF finetune stage.
@@
    Comprehensive ablations and results have demonstrated that our method can achieve competitive performance with training-based methods (e.g., DPO) in hallucination benchmarks while maintaining the general ability.
@@ 2 Related Work
Multimodal Large Language Model (MLLM). MLLMs have witnessed remarkable advancements these days. Previous arts liu2023visual ; li2022blip ; instructblip have shaped the paradigm of current MLLMs’ architecture: a vision encoder zhai2023sigmoid ; radford2021learning to process visual input, an LLM touvron2023llama ; touvron2023llama2 to reason and generate text, and a cross-modal projector li2022blip ; alayrac2022flamingo to bridge the gap between the visual and textual representations. The training for MLLMs typically involves two main stages: pre-training and post-training. The large-scale pre-training stage radford2018improving provides the model with a strong foundation of general knowledge. The post-training alignment stage consists of two phases: supervised fine-tuning (SFT) radford2018improving and reinforcement learning from human feedback (RLHF) rafailov2024direct ; stiennon2020learning ; bai2022training ; schulman2017proximal . This process refines the model’s task-specific performance and encourages alignment with human preferences. Building upon this foundation, current research continuously pushes the boundaries of their capabilities qwenvl ; hong2024cogvlm2 ; laurenccon2024building ; liu2024llavanext ; chen2024expanding ; li2024survey . Meanwhile, some research investigates alternative architectures that could shape the future of MLLMs, such as Omni fu2024vita ; li2025baichuan ; ye2024x ; zhen20243d , MoE wu2024deepseek ; lin2024moe ; li2024aria , Encoder-Free diao2025evev2 ; fuyu-8b ; chen2024single , and Any-to-Any team2024chameleon ; wang2024emu3 ; zhou2024transfusion ; chen2025janus .
@@ 3 Preliminary
Figure 2: Comparison between our method and existing methods (DPO rafailov2024direct and VCD leng2024mitigating ) in the training stage and inference stage.
@@ 4.2 Negative Samples Learning
where $\pi_{\theta}$ is the parameter of the MLLM. This forces $g_{\phi}$ to learn transformations of $v$ that align with the error distribution in $y_{l}$, effectively mapping $v$ to a “hallucination-primed” embedding space.
@@ 5.1 Experiment Setup
SEED MathVista MMStar MMMU MM-Vet MMHal Hallusion
@@
   Ours (Neg. Only) 58.60+1.23 27.8-0.7 33.00-0.30 34.7+1.1 25.1+0.7 1.80-0.17 0.70+0.05 40.38+2.31 36.59+0.72
   Ours (Pos. & Neg.) 58.55+1.18 28.0-0.5 34.53+1.23 34.5+0.9 25.0+0.6 1.77-0.20 0.69+0.04 40.48+2.41 36.84+0.97
@@
   Ours (Neg. Only) 58.60+4.12 28.3+1.7 33.20+0.20 34.4-1.2 29.4-0.3 2.00+0.39 0.66+0.02 40.17+2.32 37.34+1.13
   Ours (Pos. & Neg.) 58.61+4.13 27.9+1.3 34.47+1.47 34.1-1.5 29.5-0.2 1.66+0.05 0.60-0.04 39.54+1.69 37.35+1.14
@@
   Ours (Neg. Only) 58.57+1.14 28.7+1.9 33.07-0.06 34.3-0.6 25.6+0.1 1.70-0.20 0.72+0.06 39.85+3.89 36.68+1.06
   Ours (Pos. & Neg.) 58.56+1.13 28.4+1.6 34.53+1.40 34.0-0.9 25.5+0.0 1.86-0.04 0.69+0.03 39.43+3.47 36.73+1.11
@@
   Ours (Neg. Only) 58.62+1.75 27.5+0.6 33.20+0.93 34.4+1.4 26.1-0.5 1.83-0.35 0.69+0.01 39.75+8.20 36.60+2.07
   Ours (Pos. & Neg.) 58.59+1.72 28.1+1.2 34.61+2.34 34.1+1.1 27.3+0.7 1.80-0.38 0.70+0.02 39.96+8.41 37.11+2.58
@@
Implementation Details. We conduct our experiments on LLaVA 1.5-7B liu2024improved , training only the image projection layer while keeping all other parameters frozen. For training, we use the above four hallucination-related preference datasets: RLHF-V yu2024rlhf is trained for 2 epochs, while the remaining datasets are trained for 1 epoch each. Hyperparameters for contrastive decoding follow the configuration recommended in VCD leng2024mitigating , ensuring consistency with this baseline approach. For the DPO baseline, we follow the training setting of BPO pi2024strengthening .
@@ 5.2 Quantitative Results
Comparison to VCD. While VCD marginally improves POPE accuracy, it degrades performance on complex benchmarks like MathVista (--0.9 %) and open-end benchmarks like HallusionBench (--0.2 %). Our method outperforms VCD across all metrics, demonstrating that learned negative embeddings better capture authentic hallucination patterns than static noise perturbations.
SEED MM-Vet Hallusion POPE
@@
Table 3: Ablation study of the type of negative image embedding used to contrastive decoding. “Add Noise” is adding noise to the image to get negative image embedding which is adopted by VCD leng2024mitigating , “Other image” means randomly sampling another image as negative image embedding, and “Nega Projection” is our method trained on BPO pi2024strengthening which utilizes a negative image projection to get negative image embedding. For POPE pope , we report the results of the adversarial set here.
SEED MM-Vet Hallusion POPE
@@ 5.3 Ablation Studies
SEED MM-Vet Hallusion POPE
@@
5.4 Qualitative Analysis
Case Study. In the 1st row of Figure 3, VCD fails to address the hallucination issue in the table scene, whereas both DPO and our method succeed. However, on the right side, DPO provides an incorrect answer and responds oddly by self-questioning and using another language (e.g., Japanese here) due to the likelihood displacement. In the 2nd row (samples from SEED-Bench), VCD and DPO incorrectly answered general ability questions that the baseline model (LLaVA-1.5 7B) originally answered correctly, while, our method can preserve baseline model’s original capability.

### v2

@@ Abstract
Although multimodal large language models (MLLMs) exhibit remarkable reasoning capabilities on complex multimodal understanding tasks, they still suffer from the notorious “hallucination” issue: generating outputs misaligned with obvious visual or factual evidence. Currently, training-based solutions, like direct preference optimization (DPO), leverage paired preference data to suppress hallucinations. However, they risk sacrificing general reasoning capabilities due to the likelihood displacement. Meanwhile, training-free solutions, like contrastive decoding, achieve this goal by subtracting the estimated hallucination pattern from a distorted input. Yet, these handcrafted perturbations (e.g., add noise to images) may poorly capture authentic hallucination patterns. To avoid these weaknesses of existing methods, and realize “robust” hallucination mitigation (i.e., maintaining general reasoning performance), we propose a novel framework: Decoupling Contrastive Decoding (DCD). Specifically, DCD decouples the learning of positive and negative samples in preference datasets, and trains separate positive and negative image projections within the MLLM. The negative projection implicitly models real hallucination patterns, which enables vision-aware negative images in the contrastive decoding inference stage. Our DCD alleviates likelihood displacement by avoiding pairwise optimization and generalizes robustly without handcrafted degradation. Extensive ablations across hallucination benchmarks and general reasoning tasks demonstrate the effectiveness of DCD, i.e., it matches DPO’s hallucination suppression while preserving general capabilities and outperforms the handcrafted contrastive decoding methods. Code is available in https://github.com/HKUST-LongGroup/DCD.
@@ 1 Introduction
Today’s multimodal large language models (MLLMs) liu2024improved ; chen2023minigpt ; bai2023qwen ; dubey2024llama ; chen2024expanding have demonstrated remarkable general reasoning capabilities by integrating visual and textual understanding, facilitating applications such as medical image analysis li2024llava ; liu2023medical and multimodal search engines caffagni2024wiki . Despite their versatility, a critical limitation persists: MLLMs may generate outputs that contradict obvious factual evidence or misrepresent visual inputs, known as the hallucination problem bai2024hallucination ; liu2024survey ; yu2024hallucidoctor ; zhou2023analyzing . For instance, models may describe objects absent from an image (e.g., claiming a “dog” in a cat-only scene) or fabricate implausible relationships (e.g., asserting “a person riding a bicycle” when only a bicycle is present). Such hallucinations erode users trust and hinder deployment in high-stakes domains like healthcare li2024llava or autonomous driving cui2024survey .
Figure 1: Comparison between existing hallucination mitigation methods and DCD. (a) Training-based method (e.g., DPO rafailov2024direct ): DPO directly optimizes the likelihood gap between positive (correct) and negative (hallucinatory) responses using preference datasets. However, maximizing this gap ($y^{+}$ vs. $y^{-}$) can inadvertently lower the probability of both responses, causing likelihood displacement and potential degradation of general reasoning capabilities. Here, $v$, $x$, $y^{+}$, and $y^{-}$ denote images, questions, positive responses, and negative responses, respectively; $\theta$ represents model parameters, and $\alpha$ is the contrastive decoding coefficient. (b) Training-free method (e.g., VCD leng2024mitigating ) vs. DCD: Traditional contrastive decoding (VCD) reduces hallucinations by comparing model outputs from original ($v^{+}$) and artificially distorted ($v^{-}$; e.g., noise-added) visual inputs at inference time.
@@
    Decoupled Learning for Robust Alignment. We propose Decoupled Contrastive Decoding (DCD), the first framework to separate positive/negative sample optimization from preference datasets in MLLM training. It alleviates the likelihood displacement problem in DPO rafailov2024direct , preserving general capabilities while mitigating hallucinations.
@@
    Vision-Aware Hallucination Suppression. We introduce a learnable negative image projector trained on real hallucinatory samples. Unlike handcrafted perturbations (e.g., VCD leng2024mitigating ), this projector generates distortions grounded in actual MLLM errors, enabling precise suppression of hallucinations.
@@
    Paradigm Shift in Preference Learning. We reveal that negative samples alone suffice for hallucination mitigation, challenging the prevailing preference-learning paradigm—showing that explicit modeling of errors (not just positive alignment) is critical for robustness.
  • 4)
    Comprehensive Experiments. Extensive ablations and results demonstrate that our method achieves competitive performance with training-based methods (e.g., DPO rafailov2024direct ) on hallucination benchmarks while maintaining general ability.
@@ 2 Related Work
Multimodal Large Language Model (MLLM). MLLMs have witnessed remarkable advancements these days. Previous arts liu2023visual ; li2022blip ; instructblip ; li2023zero have shaped the paradigm of current MLLMs’ architecture: a vision encoder zhai2023sigmoid ; radford2021learning to process visual input, an LLM touvron2023llama ; touvron2023llama2 to reason and generate text, and a cross-modal projector li2022blip ; alayrac2022flamingo ; lei2024seeing to bridge the gap between the visual and textual representations. The training for MLLMs typically involves two main stages: pre-training and post-training. The large-scale pre-training stage radford2018improving provides the model with a strong foundation of general knowledge. The post-training alignment stage consists of two phases: supervised fine-tuning (SFT) radford2018improving and reinforcement learning from human feedback (RLHF) rafailov2024direct ; stiennon2020learning ; bai2022training ; schulman2017proximal . This process refines the model’s task-specific performance and encourages alignment with human preferences. Building upon this foundation, current research continuously pushes the boundaries of their capabilities qwenvl ; hong2024cogvlm2 ; laurenccon2024building ; liu2024llavanext ; chen2024expanding ; li2024survey . Meanwhile, some research investigates alternative architectures that could shape the future of MLLMs, such as Omni fu2024vita ; li2025baichuan ; ye2024x ; zhen20243d , MoE wu2024deepseek ; lin2024moe ; li2024aria , Encoder-Free diao2025evev2 ; fuyu-8b ; chen2024single , and Any-to-Any team2024chameleon ; wang2024emu3 ; zhou2024transfusion ; chen2025janus .
@@ 3 Preliminary
Figure 2: Comparison of DCD with DPO rafailov2024direct and VCD leng2024mitigating in the training and inference stages. (a) Training stage: DPO jointly optimizes positive–negative responses, risking likelihood displacement. Our method (DCD) separately learns positive and negative image projections to avoid this issue. (b) Inference stage: VCD uses artificial noise as negative inputs, whereas DCD leverages learned negative visual features that reflect authentic hallucination patterns, enhancing effective hallucination suppression.
@@ 4.2 Negative Samples Learning
where $\theta$ is the parameter of the MLLM. This forces $g_{\phi}$ to learn transformations of $v$ that align with the error distribution in $y_{l}$, effectively mapping $v$ to a “hallucination-primed” embedding space.
@@ 5.1 Experiment Setup
SEED MathVista MMStar MMMU $M$-Vet MMHal Hallusion
@@
  SimPO meng2024simpo 58.00 28.1 33.40 33.0 26.7 1.95 0.69 36.70 35.98
  Ours (Neg. Only) 58.60+0.60 27.8-0.7 33.00-0.40 34.7+1.1 25.1-1.6 1.80-0.17 0.70+0.05 40.38+2.31 36.59+0.61
  Ours (Pos. & Neg.) 58.55+0.55 28.0-0.5 34.53+1.13 34.5+0.9 25.0-1.7 1.77-0.20 0.69+0.04 40.48+2.41 36.84+0.86
@@
  SimPO meng2024simpo 57.07 27.6 32.47 34.3 27.3 1.24 0.80 39.53 36.58
  Ours (Neg. Only) 58.60+1.53 28.3+0.7 33.20+0.20 34.4-1.2 29.4-0.3 2.00+0.39 0.66+0.02 40.17+0.64 37.34+0.76
  Ours (Pos. & Neg.) 58.61+1.54 27.9+0.3 34.47+1.47 34.1-1.5 29.5-0.2 1.66+0.05 0.60-0.04 39.54+0.01 37.35+0.77
@@
  SimPO meng2024simpo 57.89 27.8 32.80 33.2 27.1 1.67 0.71 36.80 36.24
  Ours (Neg. Only) 58.57+0.68 28.7+0.9 33.07-0.06 34.3-0.6 25.6-1.5 1.70-0.20 0.72+0.06 39.85+3.05 36.68+0.44
  Ours (Pos. & Neg.) 58.56+0.67 28.4+0.6 34.53+1.40 34.0-0.9 25.5-1.6 1.86-0.04 0.69+0.03 39.43+2.63 36.73+0.49
@@
  SimPO meng2024simpo 58.24 28.0 31.47 32.7 27.0 1.74 0.75 30.28 34.98
  Ours (Neg. Only) 58.62+0.38 27.5+0.6 33.20+0.93 34.4+1.4 26.1-0.9 1.83-0.35 0.69+0.01 39.75+8.20 36.60+1.62
  Ours (Pos. & Neg.) 58.59+0.35 28.1+1.2 34.61+2.34 34.1+1.1 27.3+0.3 1.80-0.38 0.70+0.02 39.96+8.41 37.11+2.13
@@
Implementation Details. We conduct our experiments on LLaVA 1.5-7B liu2024improved , training only the image projection layer while keeping all other parameters frozen. For training, we use the above four hallucination-related preference datasets: RLHF-V yu2024rlhf is trained for 2 epochs, while the remaining datasets are trained for 1 epoch each on NVIDIA A100 80GB. Hyperparameters for contrastive decoding follow the configuration recommended in VCD leng2024mitigating , ensuring consistency with this baseline approach. For the DPO baseline, we follow the training setting of BPO pi2024strengthening .
@@ 5.2 Quantitative Results
Comparison to VCD. While VCD marginally improves POPE accuracy, it degrades performance on complex benchmarks like MathVista (-0.9 %) and open-end benchmarks like HallusionBench (-0.2 %). Our method outperforms VCD across all metrics, demonstrating that learned negative embeddings better capture authentic hallucination patterns than static noise perturbations.
SEED $M$-Vet Hallusion POPE
@@
Table 3: Ablation study of the type of negative image embedding used to contrastive decoding. “Add Noise” is adding noise to the image to get negative image embedding which is adopted by VCD leng2024mitigating , “Other image” means randomly sampling another image as negative image embedding, and “Nega Projection” is our method trained on BPO pi2024strengthening which utilizes a negative image projection to get negative image embedding. We present the adversarial set results for POPE pope .
SEED $M$-Vet Hallusion POPE
@@ 5.3 Ablation Studies
SEED $M$-Vet Hallusion POPE
@@
5.4 Qualitative Analysis
Case Study. In the 1st row of Figure 3, VCD fails to address the hallucination issue in the table scene, whereas both DPO and our method succeed. However, on the right side, DPO provides an incorrect answer and responds oddly by self-questioning and using another language (e.g., Japanese here) due to the likelihood displacement. In the 2nd row (samples from SEED-Bench), VCD and DPO incorrectly answered general ability questions that the baseline model (LLaVA-1.5 7B) originally answered correctly, while our method can preserve baseline model’s original capability.
@@ 6 Conclusion
Acknowledgment
This work was supported by the National Natural Science Foundation of China Young Scholar Fund (62402408) and the Hong Kong SAR RGC Early Career Scheme (26208924).
Appendix
Appendix A Social Impacts
Our work addresses the critical challenge of hallucination in multimodal large language models (MLLMs), with profound implications for the safe deployment of AI systems across socially sensitive domains. By mitigating factual contradictions and visual misrepresentations, our method enhances model reliability in high-stakes applications such as medical diagnostics and autonomous driving. Beyond safety-critical scenarios, our approach fosters trust in AI-assisted decision-making tools for education, legal documentation, and content moderation by ensuring outputs align with observable evidence. This reliability is particularly crucial for combating misinformation in an era of AI-generated content proliferation. Additionally, our findings on the importance of negative sample learning offer insights for developing more efficient alignment frameworks, potentially democratizing access to robust MLLMs for resource-constrained institutions.
Appendix B Limitations
While our method achieves robust hallucination mitigation, two key limitations warrant consideration. First, the contrastive decoding framework inherently doubles computational overhead during inference due to parallel processing of original and negative-projected image features. Second, our negative image projection relies on the quality and diversity of hallucination patterns in preference datasets. While current datasets predominantly cover common object hallucinations (e.g., spurious object mentions), they may underrepresent complex multimodal hallucinations involving spatial reasoning or causal relationships. Future work could explore adaptive weighting mechanisms to handle such edge cases.
Appendix C Additional Experiments
Results on Qwen 2.5 VL 3B. As shown in Table A1 and Table A2, on a lightweight yet strong backbone, DCD preserves—or slightly improves—general capability while consistently suppressing hallucinations. Averaged over settings, our method surpasses DPO in three of four preference regimes, with ”Pos.&Neg.” typically the most stable variant. Overall, these appendix results support DCD’s supervision- and model-agnostic transferability: learned negative embeddings—especially with joint positive-negative training—capture authentic hallucination patterns and extend cleanly to compact backbones, pointing to straightforward scaling on Qwen variants and broader preference signals.
Appendix D Theoretical Foundation
1. Key Proposition: A Sufficient Condition for Eliminating Likelihood Displacement
From Pairwise to Decoupled Optimization. DPO maximizes a paired log gap:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,v,y^{+},y^{-})}\!\left[\log\sigma\!\left(\beta\left[\log\pi_{\theta}(y^{+}\mid x,v)-\log\pi_{\theta}(y^{-}\mid x,v)\right]\right)\right],$
which guarantees the gap widens but is agnostic to the absolute values of $\log\pi_{\theta}(y^{+}\mid x,v)$ and $\log\pi_{\theta}(y^{-}\mid x,v)$. As a result, both likelihoods can drift downward—the likelihood displacement effect that degrades general reasoning (see Figure 1(a) DPO in the paper).
Our Decoupled Objective. Instead, we minimize two independent cross-entropies:
$\min_{\psi}\;\mathbb{E}_{(x,v,y^{+})}\!\left[-\log\pi_{\theta}\!\left(y^{+}\mid x,g_{\psi}(v)\right)\right],\qquad\min_{\phi}\;\mathbb{E}_{(x,v,y^{-})}\!\left[-\log\pi_{\theta}\!\left(y^{-}\mid x,g_{\phi}(v)\right)\right],$
and combine the results only at inference time:
$\widehat{\operatorname{logit}}\;=\;(1+\alpha)\,\operatorname{logit}_{\psi}\;-\;\alpha\,\operatorname{logit}_{\phi}.$
As shown in Figure 1(a) Ours, this ensures that $\log\pi_{\theta}(y^{+}\mid x,v)$ increases while $\log\pi_{\theta}(y^{-}\mid x,v)$ is suppressed. This construction provides a lower-bound guarantee on general reasoning performance, consistent with our empirical results (Table 1, SEED-Bench).
2. Necessity and Sufficiency of the Negative Projector
Theoretical View. We interpret the negative image projection $g_{\phi}$ as learning an adversarial negative distribution $q_{\phi}(v)$ in the vision-feature space $\mathcal{V}$, which maximizes
$\mathrm{KL}\!\left(p(y\mid x,v)\;\Big\|\;p\!\left(y\mid x,g_{\phi}(v)\right)\right),$
equivalent to maximizing the InfoNCE lower bound. Thus, learning only the negative projector implicitly provides a gradient-shaped penalty during inference, which explains why our Neg-only training consistently reduces hallucination.
Empirical Evidence. As shown in Table 1, the Neg-only variant nearly matches Pos + Neg on hallucination benchmarks, and clearly outperforms Pos-only. To our knowledge, ours is the first work to show that negative-only learning can suffice.
General Performance Hallucination Average
SEED MathVista MMStar MMMU $M$-Vet MMHal Hallusion
Score Rate $\downarrow$
  Qwen 2.5 VL 3B 66.67 61.0 53.27 44.4 52.5 2.01 0.69 50.68 54.75
  + VCD leng2024mitigating 65.15 64.1 52.53 44.9 52.5 1.96 0.71 49.21 54.73
Fine-tuned on RLHF-V yu2024rlhf
  DPO rafailov2024direct 66.61 60.9 53.53 44.8 51.3 1.83 0.73 50.81 54.66
  SimPO meng2024simpo 66.89 61.2 53.07 44.9 50.1 2.07 0.68 49.63 54.30
  Ours (Neg. Only) 65.00-1.89 62.4+1.20 53.53+0.00 44.3-0.60 53.2+1.90 1.98-0.09 0.70+0.02 51.94+1.13 55.06+0.40
  Ours (Pos. & Neg.) 66.30-0.59 62.7+1.50 54.80+1.27 44.4-0.50 54.5+3.20 2.43+0.36 0.57-0.11 52.37+1.56 55.85+1.19
Fine-tuned on BPO pi2024strengthening
  DPO rafailov2024direct 66.96 63.8 53.13 45.3 52.4 1.84 0.69 52.66 55.71
  SimPO meng2024simpo 66.92 63.2 53.67 45.1 51.5 2.35 0.55 53.68 55.68
  Ours (Neg. Only) 67.11+0.15 60.2-3.60 52.93-0.74 46.6+1.30 53.8+1.40 1.97-0.38 0.68+0.13 53.31-0.37 55.66-0.05
  Ours (Pos. & Neg.) 67.12+0.16 63.2-0.60 53.07-0.60 46.9+1.60 53.2+0.80 2.55+0.20 0.52-0.03 53.26-0.42 56.13+0.42
Fine-tuned on RLAIF-V yu2024rlaif
  DPO rafailov2024direct 66.84 60.8 53.13 44.6 52.9 1.58 0.76 50.16 54.74
  SimPO meng2024simpo 66.32 60.2 52.67 45.1 45.6 2.55 0.59 46.37 52.71
  Ours (Neg. Only) 64.70-2.14 61.1+0.30 52.87-0.26 46.7+1.60 53.0+0.10 1.99-0.41 0.71-0.05 52.47+2.31 55.14+0.40
  Ours (Pos. & Neg.) 65.61-1.23 60.9+0.10 50.27-2.86 45.8+0.70 53.5+0.60 2.30+0.72 0.61-0.15 52.68+2.52 54.79+0.06
Fine-tuned on VLFeedback li2024vlfeedback
  DPO rafailov2024direct 66.74 60.5 52.33 44.6 53.1 2.14 0.66 48.27 54.26
  SimPO meng2024simpo 66.96 62.3 52.87 44.6 51.9 2.33 0.68 48.05 54.45
  Ours (Neg. Only) 65.80-1.16 61.2-1.10 53.13+0.26 45.2+0.60 53.0-0.10 2.21-0.12 0.66-0.00 51.79+3.52 55.02+0.57
  Ours (Pos. & Neg.) 66.25-0.71 61.0-1.30 53.60+0.73 45.6+1.00 54.7+1.60 3.15+0.82 0.46-0.20 51.69+3.42 55.47+1.03
Table A1: Performance comparison on general and hallucination benchmarks for Qwen 2.5 VL 3B. ‘Neg. Only’ means only trained on negative samples of preference datasets, ‘Pos. & Neg.’ is trained with both positive and negative samples, $\downarrow$ indicates lower is better, and denotes that the MMHal values are not counted in the Average score. $\dagger$ For benchmarks requiring GPT evaluation, we follow the same setting as the main table (e.g., GPT-4o 24-05-13).
Random Popular Adversarial
Acc F1 Acc F1 Acc F1
Woodpeckeryin2024woodpecker 85.51 86.67 83.51 84.33 82.35 83.00
Qwen 2.5 VL 3B 88.90 87.67 87.87 86.68 86.67 85.55
+ VCD leng2024mitigating 88.93 87.79 87.33 86.27 85.80 84.85
Fine-tuned on RLHF-V yu2024rlhf
DPO rafailov2024direct 88.83 87.58 87.83 86.61 86.60 85.46
SimPO meng2024simpo 88.07 86.59 87.27 85.82 86.23 84.84
Ours (Neg. Only) 89.57 88.54 88.30 87.32 86.80 85.92
Ours (Pos. & Neg.) 89.30 88.18 87.93 86.87 86.87 85.87
Fine-tuned on BPO pi2024strengthening
DPO rafailov2024direct 89.40 88.29 88.20 87.13 86.87 85.89
SimPO meng2024simpo 88.17 86.72 87.23 85.82 86.07 84.72
Ours (Neg. Only) 91.37 90.77 89.07 88.58 86.76 86.51
Ours (Pos. & Neg.) 90.90 90.20 89.33 88.70 86.63 86.23
Fine-tuned on RLAIF-V yu2024rlaif
DPO rafailov2024direct 89.23 88.08 88.10 86.99 86.67 85.64
SimPO meng2024simpo 89.57 88.46 88.40 87.33 86.90 85.93
Ours (Neg. Only) 91.37 90.80 88.40 88.02 86.07 85.95
Ours (Pos. & Neg.) 90.30 89.56 88.53 87.89 85.80 85.41
Fine-tuned on VLFeedback li2024vlfeedback
DPO rafailov2024direct 87.77 86.21 87.07 85.53 86.00 84.52
SimPO meng2024simpo 87.03 85.20 86.40 84.59 85.50 83.74
Ours (Neg. Only) 90.03 89.08 88.47 87.58 87.07 86.28
Ours (Pos. & Neg.) 89.27 88.17 88.00 86.96 85.90 85.02
Table A2: Performance comparison on POPE pope with Qwen 2.5 VL 3B. “Neg. Only” uses only negative samples from preference datasets; “Pos. & Neg.” uses both positive and negative samples.